The combined age of a mother and her twin daughters is 58. The mother was 25 years old when the twins where born. Write and solve an equation to find the age of each of the three people.
Answer by mananth(16946)

let the daughter's age be x
mother's age = y
2x+y =58
y-x=25
multiply equation by 2
2y-2x=50
add to the first equation
2x+y+2y-2x=50+58
3y=108
/3
y=36
mother's age =36
2x+y=58
2x+36=58
2x=58-36
2x=22
x=11
Daughter's ages = 11 years